A non-inverting amplifier is working on 120 mV with input resistance of 300 and feedback resistance of 1500. Calculate its gain and design the circuit by showing all parameters.

1
Expert's answer
2022-01-10T17:01:03-0500

feedback resistance (R1)= 1500 ohm


input resistance (R2)= 300 ohm


Gain=1+R1/R2

= 1+1500/300=6